You buy items costing $

2300

and finance the cost with a fixed installment loan for

24

months at

5

% simple interest per year.

a.

What is the

finance charge?

b.

What

is

your monthly payment


1
Expert's answer
2021-08-30T05:21:19-0400

a) finance charge= "\frac{PRT}{100}=\frac{2300\times 5\times 24/12}{100}=\$230" (for 24months)


b) monthly payment "=\frac{2300}{24months}+\frac{230}{24months}=95.83+9.583=\$105.413"